OpenStudy (anonymous):

Evaluate the integral. (x-3)/x dx from 1 to 4

OpenStudy (campbell_st):

rewrite the problem by splitting the numerator \[\int\limits_{1}^{4} \frac{x}{x} - \frac{3}{x} dx = \int\limits_{1}^{4} 1 - \frac{3}{x} dx\]

OpenStudy (anonymous):

For the integrand substitute u=x-3 and du=dx \[\int\limits_{1}^{4}\frac{ u }{ u+3 }du\] Do long division: \[\int\limits\limits_{1}^{4}1-\frac{3}{ u+3 }du\] Integrate term by term: \[\int\limits_{1}^{4}1du-3\int\limits_{1}^{4}\frac{ 1 }{ u+3 }du\] Substitute s=u+3 and ds=du \[\int\limits_{1}^{4}1du-3\int\limits_{1}^{4}1/sds\] Integral of 1/s is log(s) so we get u-3log(s) sub back in for u (x-3)log(x)-3 By plugging in the limits of integration we get: 3-3log(4)

OpenStudy (campbell_st):

its not that difficult a question ... it doesn't need substitution